40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 173     Entry time: Tue Oct 22 17:17:59 2013
Author: Koji 
Type: General 
Category: General 
Subject: H1OMC cabling 

[Chub, Jeff, Koji]

We worked on the wiring and routing of the cables.

- The cables for the PZT was installed first.

- Pins for the mighty mouse connector were crimped on the PZT wires

- Checked the wiring diagram (D1300589) to find the pinouts.
  Pin1 of the mighty mouse is connected to PZT2+, Pin2 to PZT2-, PIn3 to PZT1+, and Pin4 to PZT1-

- Then QPD and PD cables are fixed on the cable harness.

- The QPD/PD cables are attached on the diode housings.

During this process one of the DCPD mounts moved. The fixing screws were not torqued enough.
This means that all of the FC layers need to be removed and the DCPD housing should be aligned again.

- We continued on the cabling. The cables were routed on the top (cable) side.

- Some of the cable pegs were tightened by PEEK cable ties.

- We found that Pin1 and Pin2 of the PZT cables were not intact anymore.

- We ask Chub to work on the PZT pins tomorrow. The PD alignment will be taken tonight or tomorrow.

ELOG V3.1.3-